# Towards user-centered interactive medical image segmentation in VR with an assistive AI agent

### v2

@@ Abstract
Crucial in disease analysis and surgical planning, manual segmentation of volumetric medical scans (e.g. MRI, CT) is laborious, error-prone, and challenging to master, while fully automatic algorithms can benefit from user feedback. Therefore, with the complementary power of the latest radiological AI foundation models and virtual reality (VR)’s intuitive data interaction, we propose SAMIRA, a novel conversational AI agent that assists users with localizing, segmenting, and visualizing 3D medical concepts in VR. Through speech-based interaction, the agent helps users understand radiological features, locate clinical targets, and generate segmentation masks that can be refined with just a few point prompts. The system also supports true-to-scale 3D visualization of segmented pathology to enhance patient-specific anatomical understanding. Furthermore, to determine the optimal interaction paradigm under near-far attention-switching for refining segmentation masks in an immersive, human-in-the-loop workflow, we compare VR controller pointing, head pointing, and eye tracking as input modes. With a user study, evaluations demonstrated a high usability score (SUS=90.0±9.0), low overall task load, as well as strong support for the proposed VR system’s guidance, training potential, and integration of AI in radiological segmentation tasks.
Index Terms:
@@ III System Overview
IV Methods and Materials
IV-A Interaction Paradigm Evaluation
@@
IV-B SAMIRA - User Interface and Workflow
The full workflow of our interactive medical image segmentation leverages SAMIRA, our assistive AI agent via the VR user interface described in Section III. The interface enables SAMIRA to operate on unseen cases without ground truth segmentations, and supplies reference images and case-relevant descriptive guidance in real time. This supports the potential for future clinical and educational applications beyond the controlled evaluation settings showcased in our study.
@@ IV-B SAMIRA - User Interface and Workflow
IV-C SAMIRA - Segmentation Algorithm and Retrieval-Augmented Generation
@@
IV-C1 Interactive segmentation with foundation models
@@
While point-prompt-based revision with SAM2 helps ensure the accuracy of the segmentation, it is desirable to keep the needs at minimum for the efficiency of the workflow. Thus, to gauge the baseline performance of our proposed segmentation method in the absence of prompt-based revision, we conducted a standalone ablation study using the Pretreat-MetsToBrain-Masks [24] and LiTS [25] datasets, by comparing the accuracy of a “BiomedParse seeding + SAM2 propagation with IoU-based early stopping” pipeline versus a “BiomedParse seeding + original SAM2 propagation” one. To accomplish this, we randomly selected 40 volumes (20 brain tumor MRI scans and 20 liver tumor CT scans) from the datasets, excluding cases used in the full workflow study. For each volume, a slice with visible tumor was manually selected and submitted to the inference server along with a natural language prompt —“show me the brain tumor” for brain cases and “show me the liver tumor” for liver cases. Here, BiomedParse generates an initial tumor mask for the selected slice, which was then propagated bi-directionally (in the superior and inferior directions) using SAM2. Using the paired sample Wilcoxon signed-rank test, results indicate that the break condition significantly improved accuracy for liver tumor CT scans ($p=0.0024<0.05$), while having a smaller but still significant effect for brain tumor MRIs ($p=0.0039<0.05$), as summarized in Table I. The stronger effect observed in CT scans may be attributed to the higher levels of noise and lower soft-tissue contrast, which can cause SAM2, originally trained on natural images, to mistake noise for anatomical structures. The break condition prevents propagation of spurious segmentations across slices, which is particularly helpful in noisier CTs. Compared to recent SAM2-based methods for 3D medical segmentation [11], our approach introduces both a language-based initialization and a propagation stopping rule, enhancing accuracy and reducing user burden. These findings support the inclusion of the IoU threshold as an essential mechanism for robust slice-wise mask propagation in clinical datasets.
TABLE I: Mean 3D Dice scores for automatic SAM2 propagation with and without the IoU break condition.
Modality Target IoU Break 3D Dice (%, Mean $\pm$ Std)
@@
IV-C2 RAG-based guidance system
@@
IV-D User Study and Evaluation Metrics
@@
IV-D1 User-Interaction Paradigm Evaluation
@@
IV-D2 Full Workflow Study of SAMIRA
@@ IV-D3 Statistical Analysis
V Results
V-A Interaction Paradigm Evaluation
TABLE II: Comparison of interaction paradigms for segmentation refinement. Values are shown as mean $\pm$ standard deviation. The best score is in bold fonts. NASA-TLX is out of 100.
@@ V-A Interaction Paradigm Evaluation
Controller $\textbf{99.25}\pm\textbf{00.25}$ $\textbf{220.3}\pm\textbf{79.3}$ $18.8\pm 14.5$ $\textbf{0.51}\pm\textbf{1.91}$
Head Pointing $99.21\pm 0.30$ $248.8\pm 78.5$ $\textbf{16.8}\pm\textbf{13.9}$ $0.20\pm 1.56$
Eye Tracking $99.13\pm 0.46$ $251.1\pm 78.7$ $26.6\pm 15.4$ $-0.71\pm 1.77$
V-A1 Accuracy, Completion Time, and NASA-TLX
The metrics for evaluating the three interaction paradigms are shown in Table  II, with non-significant differences in overall scores between groups ($p>0.05$). For 3D Dice (segmentation accuracy), all paradigms yielded highly accurate refined masks, significantly above the starting 3D Dice score of 0.91 ($p<0.05$). The Controller paradigm had the highest accuracy score with the least variability ($99.25\pm 0.25$%), followed by Head Pointing ($99.21\pm 0.30$%), then Eye Tracking ($99.13\pm 0.46$%). For completion time, the ranking remained the same: Controller was the fastest ($220.3\pm 79.3s$) followed by Head Pointing ($248.8\pm 78.5s$) and Eye Tracking ($251.1\pm 78.7s$). The overall NASA-TLX scores (out of 100) slightly favored Head Pointing, which had the lowest overall task load ($16.8\pm 13.9$), followed by Controller ($18.8\pm 14.5$) and Eye Tracking ($26.6\pm 15.4$). While overall task load did not demonstrate significant differences, one sub-item showed a significant effect: mental demand was significantly lower for Head Pointing ($16.7\pm 22.5$) compared to Controller ($30.0\pm 23.9$, $p=0.0123$) and Eye Tracking ($38.7\pm 25.5$, $p=0.0329$), as illustrated in Figure 6. No other NASA-TLX sub-items’ differences reached statistical significance.
@@ V-A2 Composite Scores
V-A3 Point Prompt Efficiency
@@ V-A4 User preferences
V-B Full Segmentation Workflow with SAMIRA
V-B1 Segmentation accuracy
In terms of segmentation accuracy (3D Dice), the workflow yielded high scores. For the brain MRI data, refined masks (94.92 $\pm$ 0.52%) showed significantly higher accuracy than the user’s propagated unrefined masks (90.53 $\pm$ 10.51%, $p<0.0001$). For the liver CT data, Dice scores were comparable between unrefined (95.47 $\pm$ 0.39) and refined (95.46 $\pm$ 0.40) masks, with no significant difference, ultimately indicating high starting accuracies are preserved (Table III). This high starting accuracy is likely attributed to the liver tumor case being larger and more spherical shaped than the brain tumor case that appears as a mass centrally, but splits into multiple lobes in the superior and inferior regions. The lower difficulty of the liver tumor case is further reflected in the shorter completion time (liver: $279.6\pm 109.4$s vs. brain: $361.7\pm 144.3$ s) and fewer point prompts placed (liver: 8.8 $\pm$ 6.5 vs. brain: 27.7 $\pm$ 19.9 ). The brain tumor case likely demanded more user input due to its aforementioned irregular shape.
V-B2 Semi-quantitative questionnaire results
@@
TABLE III: 3D Dice scores before and after refinement. Asterisks denote statistically significant changes.
@@
V-B3 Qualitative feedback results

### v3

@@ Abstract
Crucial in disease analysis and surgical planning, manual segmentation of volumetric medical scans (e.g. MRI, CT) is laborious, error-prone, and challenging to master, while fully automatic algorithms can benefit from user feedback. Therefore, with the complementary power of the latest radiological AI foundation models and virtual reality (VR)’s intuitive data interaction, we propose SAMIRA, a novel conversational AI agent for medical VR that assists users with localizing, segmenting, and visualizing 3D medical concepts. Through speech-based interaction, the agent helps users understand radiological features, locate clinical targets, and generate segmentation masks that can be refined with just a few point prompts. The system also supports true-to-scale 3D visualization of segmented pathology to enhance patient-specific anatomical understanding. Furthermore, to determine the optimal interaction paradigm under near-far attention-switching for refining segmentation masks in an immersive, human-in-the-loop workflow, we compare VR controller pointing, head pointing, and eye tracking as input modes. With a user study, evaluations demonstrated a high usability score (SUS=90.0±9.0), low overall task load, as well as strong support for the proposed VR system’s guidance, training potential, and integration of AI in radiological segmentation tasks.
keywords:
@@ 3 System Overview
4 Methods and Materials
4.1 Interaction Paradigm Evaluation
@@
4.2 SAMIRA - User Interface and Workflow
The full workflow of our interactive medical image segmentation leverages SAMIRA, our assistive AI agent via the VR user interface described in Section 3. The interface enables SAMIRA to operate on unseen cases without ground truth segmentations, and supplies reference images and case-relevant descriptive guidance in real time. This supports the potential for future clinical and educational applications beyond the controlled evaluation settings showcased in our study.
@@ 4.2 SAMIRA - User Interface and Workflow
4.3 SAMIRA - Segmentation and RAG
@@
4.3.1 Interactive segmentation with foundation models
@@
While point-prompt-based revision with SAM2 helps ensure the accuracy of the segmentation, it is desirable to keep the needs at minimum for the efficiency of the workflow. Thus, to gauge the baseline performance of our proposed segmentation method in the absence of prompt-based revision, we conducted a standalone ablation study using the Pretreat-MetsToBrain-Masks [24] and LiTS [25] datasets, by comparing the accuracy of a “BiomedParse seeding + SAM2 propagation with IoU-based early stopping” pipeline versus a “BiomedParse seeding + original SAM2 propagation” one. To accomplish this, we randomly selected 40 volumes (20 brain tumor MRI scans and 20 liver tumor CT scans) from the datasets, excluding cases used in the full workflow study. For each volume, a slice with visible tumor was manually selected and submitted to the inference server along with a natural language prompt —“show me the brain tumor” for brain cases and “show me the liver tumor” for liver cases. Here, BiomedParse generates an initial tumor mask for the selected slice, which was then propagated bi-directionally (in the superior and inferior directions) using SAM2. Using the paired sample Wilcoxon signed-rank test, results indicate that the break condition significantly improved accuracy for liver tumor CT scans ($p=0.0024<0.05$), while having a smaller but still significant effect for brain tumor MRIs ($p=0.0039<0.05$), as summarized in Table 1. The stronger effect observed in CT scans may be attributed to the higher levels of noise and lower soft-tissue contrast, which can cause SAM2, originally trained on natural images, to mistake noise for anatomical structures. The break condition prevents propagation of spurious segmentations across slices, which is particularly helpful in noisier CTs. Compared to recent SAM2-based methods for 3D medical segmentation [11], our approach introduces both a language-based initialization and a propagation stopping rule, enhancing accuracy and reducing user burden. These findings support the inclusion of the IoU threshold as an essential mechanism for robust slice-wise mask propagation in clinical datasets.
Table 1: 3D Dice scores (mean±$\pm$±std) for automatic SAM2 propagation with and without the IoU break condition.
Mod. Target IoU Break 3D Dice
@@
4.3.2 RAG-based guidance system
@@
4.4 User Study and Evaluation Metrics
@@
4.4.1 User-Interaction Paradigm Evaluation
@@
4.4.2 Full Workflow Study of SAMIRA
@@ 4.4.3 Statistical Analysis
5 Results
5.1 Interaction Paradigm Evaluation
Table 2: Comparison of interaction paradigms for segmentation refinement. Values are mean $\pm$ standard deviation. The best score is in bold. NASA-TLX is out of 100.
@@ 5.1 Interaction Paradigm Evaluation
Controller 99.25 $\pm$ 0.25 220.3 $\pm$ 79.3 18.8 $\pm$ 14.5 0.51 $\pm$ 1.91
Head Pointing 99.21 $\pm$ 0.30 248.8 $\pm$ 78.5 16.8 $\pm$ 13.9 0.20 $\pm$ 1.56
Eye Tracking 99.13 $\pm$ 0.46 251.1 $\pm$ 78.7 26.6 $\pm$ 15.4 –0.71 $\pm$ 1.77
5.1.1 Accuracy, Completion Time, and NASA-TLX
The metrics for evaluating the three interaction paradigms are shown in Table  2, with non-significant differences in overall scores between groups ($p>0.05$). For 3D Dice (segmentation accuracy), all paradigms yielded highly accurate refined masks, significantly above the starting 3D Dice score of 0.91 ($p<0.05$). The Controller paradigm had the highest accuracy score with the least variability ($99.25\pm 0.25$%), followed by Head Pointing ($99.21\pm 0.30$%), then Eye Tracking ($99.13\pm 0.46$%). For completion time, the ranking remained the same: Controller was the fastest ($220.3\pm 79.3s$) followed by Head Pointing ($248.8\pm 78.5s$) and Eye Tracking ($251.1\pm 78.7s$). The overall NASA-TLX scores (out of 100) slightly favored Head Pointing, which had the lowest overall task load ($16.8\pm 13.9$), followed by Controller ($18.8\pm 14.5$) and Eye Tracking ($26.6\pm 15.4$). While overall task load did not demonstrate significant differences, one sub-item showed a significant effect: mental demand was significantly lower for Head Pointing ($16.7\pm 22.5$) compared to Controller ($30.0\pm 23.9$, $p=0.0123$) and Eye Tracking ($38.7\pm 25.5$, $p=0.0329$), as illustrated in Figure 6. No other NASA-TLX sub-items’ differences reached statistical significance.
@@ 5.1.2 Composite Scores
5.1.3 Point Prompt Efficiency
@@ 5.1.4 User preferences
5.2 Full Segmentation Workflow with SAMIRA
5.2.1 Segmentation accuracy
In terms of segmentation accuracy (3D Dice), the workflow yielded high scores. For the brain MRI data, refined masks (94.92 $\pm$ 0.52%) showed significantly higher accuracy than the user’s propagated unrefined masks (90.53 $\pm$ 10.51%, $p<0.0001$). For the liver CT data, Dice scores were comparable between unrefined (95.47 $\pm$ 0.39) and refined (95.46 $\pm$ 0.40) masks, with no significant difference, ultimately indicating high starting accuracies are preserved (Table 3). This high starting accuracy is likely attributed to the liver tumor case being larger and more spherical shaped than the brain tumor case that appears as a mass centrally, but splits into multiple lobes in the superior and inferior regions. The lower difficulty of the liver tumor case is further reflected in the shorter completion time (liver: $279.6\pm 109.4$s vs. brain: $361.7\pm 144.3$ s) and fewer point prompts placed (liver: 8.8 $\pm$ 6.5 vs. brain: 27.7 $\pm$ 19.9 ). The brain tumor case likely demanded more user input due to its aforementioned irregular shape.
5.2.2 Semi-quantitative questionnaire results
@@
Table 3: 3D Dice scores before and after refinement. Asterisks denote statistically significant changes.
@@
5.2.3 Qualitative feedback results
@@ Acknowledgements
Declarations
Conflict of interest The authors declare no conflict of interest.
Data availability The data that support the findings of this study are available from the corresponding author upon reasonable request.